# Static or Dynamic: Towards Query-Adaptive Token Selection for Video Question Answering

### v1

@@ Abstract
Video question answering benefits from the rich information available in videos, enabling a wide range of applications. However, the large volume of tokens generated from longer videos presents significant challenges to memory efficiency and model performance. To alleviate this issue, existing works propose to compress video inputs, but usually overlooking the varying importance of static and dynamic information across different queries, leading to inefficient token usage within limited budgets. To tackle this, we propose a novel token selection strategy, explore-then-select, that adaptively adjust static and dynamic information needed based on question requirements. Our framework first explores different token allocations between static frames, which preserve spatial details, and dynamic frames, which capture temporal changes. Next, it employs a query-aware attention-based metric to select the optimal token combination without model updates. Our proposed framework is plug-and-play that can be seamlessly integrated within diverse video-language models. Extensive experiments show that our method achieves significant performance improvements (up to 5.8%) among various video question answering benchmarks.
Static or Dynamic: Towards Query-Adaptive Token Selection for Video Question Answering
@@ 1 Introduction
Video Question Answering (VideoQA) has broad applications across various fields (Mogrovejo and Solorio, 2024; Zhang et al., 2024a). Compared to text, videos provide more intuitive and dynamic information, delivering richer context and details by combining visual and temporal elements. Current research primarily leverages powerful large language models to build video-language models (VideoLMs) (Lin et al., 2023; Zhang et al., 2024b), significantly enhancing AI performance in VideoQA tasks. However, the extensive visual information in long videos leads to a dramatic increase in token counts. For instance, if one frame generates 196 tokens (Li et al., 2024a), a 5-minute video sampled at 1fps would produce nearly 60,000 tokens, posing significant challenges to memory requirements and model capabilities.
Figure 1: Different question types vary in their dependence on static and dynamic information in videos. For example, Question 2 relies on fine-grained dynamic information, while Question 1 and 3 only require static frames. The frames needed to answer the questions are highlighted with corresponding colored boxes.
@@
To achieve an effective allocation between static and dynamic information in video token compression, we propose a novel token selection strategy, Explore-then-Select, that adaptively aligns video tokens with textual queries under a limited token budget. Unlike previous approaches that rely on fixed rules, our strategy autonomously and adaptively combines static and dynamic content based on the nature of the questions (e.g., action description, event sequence, or object recognition), ensuring more precise responses to diverse queries.
Specifically, we categorize video frames into static and dynamic frames. Static frames are fully retained to preserve essential spatial details, such as objects, while dynamic frames are selectively processed, retaining only a subset of tokens to capture critical temporal changes and motion patterns. To optimize token allocation between these two types of frames, Explore-then-Select uses a two-stage process. In the exploration stage, we construct a search space consisting of various combinations of static and dynamic frames, where each combination generates a token subsequence of the same length (due to a fixed token budget). By adjusting the proportion of static and dynamic frames, we can prioritize either static details or dynamic changes based on question requirements. In the selection stage, we evaluate each combination using an query-aware metric derived from the shallow attention layers of VideoLMs. This metric quantifies the alignment between the query and video tokens, enabling us to select the optimal combination to answer the question.
@@
  • Building on the observation that questions rely differently on static and dynamic video information, we propose a novel Explore-then-Select framework to adaptively and effectively select video tokens reflecting the optimal balance of static and dynamic information under limited token budgets.
  • To address static and dynamic information needs, we design an effective search space of static-dynamic frame combinations. During the selection phase, we employ a query-aware approach, leveraging an attention-based metric to adaptively evaluate candidates and select the optimal combination for each question.
  • We conduct experiments on both long-video and short-video benchmarks, demonstrating the effectiveness of our method. Furthermore, we provide an in-depth analysis of the framework’s characteristics and behavior.
@@ 2.1 Video Language Models
Significant progress has been made in video language model research based on LLMs. These models can be primarily classified into two types: general-purpose vision-language models Team et al. (2024); Chen et al. (2024b); OpenAI (2024); Yao et al. (2024); Ye et al. (2023) and specialized video language models Lin et al. (2023); Zhang et al. (2024c); Li et al. (2024c); Zhang et al. (2025); Liu et al. (2024). In the realm of general-purpose vision-language models, LLaVA-OneVision Li et al. (2024a) unifies single-image, multi-image, and video tasks, while Qwen2-VL Wang et al. (2024) introduces dynamic resolution support and three-dimensional positional encoding for enhanced visual feature capture. In the field of specialized models, VideoChat Li et al. (2023b) focuses on achieving in-depth understanding and interaction with video content. LongVA Zhang et al. (2024b) has successfully extended the context length of language models, transferring their advantages in long-text processing to the video domain.
@@
FastV (Chen et al., 2024a)
ZipVL (He et al., 2024b)
@@
Ours
Table 1: Feature comparison with existing methods. “Pre-Input” refers to methods which reduce tokens before feeding them into large language models, while “Video-Specific” denotes methods which leverage the unique characteristics of video data.
@@ 2.2 Visual Token Compression
Some studies Bolya et al. (2022) focus on compressing visual tokens in video encoders. Among them, the RLT Choudhury et al. (2024) effectively reducing the number of tokens by replacing repeated patches in videos with a single patch. Some works Li et al. (2024b); Qian et al. (2025); Shen et al. (2024); Lan et al. (2024) introduce additional modules for visual token compression. For instance, BLIP-2 Li et al. (2023a) innovatively introduces the Q-Former module, which utilizes learnable query vectors to transform visual features into semantic representations, achieving efficient visual token compression. Furthermore, inspired by KV cache compression Zhang et al. (2023) in long-text processing, some research adopts it in visual token compression He et al. (2024b); Chen et al. (2024a); Fu et al. (2024b). These methods optimize token usage efficiency by setting thresholds based on specific metrics to prune visual tokens.
We compare existing methods in Table 1, noting that training-free approaches mainly compress tokens within the KV cache, reducing FLOPs but not addressing the issue of excessive token input to the large language model. In contrast, methods that reduce tokens in advance typically require training. This paper introduces a novel pre-input, training-free framework for more effective compression, considering query-aware static and dynamic information balancing.
@@ 3 Preliminary
  • Visual Encoding and Embedding. The sampled frames are decomposed into non-overlapping spatiotemporal patches, which are processed by a vision encoder to extract spatiotemporal features. These features are projected into the language model’s token space via a linear projector, resulting in visual token embeddings $\bm{F}\in\mathbb{R}^{T\times H\times W\times D}$, where $T$ represents the temporal resolution, $H$ and $W$ denote the spatial resolutions, and $D$ is the token embedding dimension.
  • Multimodal Processing. The visual token embeddings $\bm{F}$ are then flattened into a sequence $\bm{T}_{v}\in\mathbb{R}^{THW\times D}$, where the sequence length is $L=T\times H\times W$. The sequence $\bm{T}_{v}$, instruction embeddings $\bm{T}_{i}$, and query embeddings $\bm{T}_{q}$ are concatenated into a unified input $\bm{T}=[\bm{T}_{i},\bm{T}_{v},\bm{T}_{q}]$, where []delimited-[][$\cdot$][ ⋅ ] denotes token concatenation. Finally, VideoLM processes the unified input sequence $\bm{T}$ to generate a textual response to the question.
Figure 2: Overview of our Explore-then-Select framework for token selection. In the exploration stage, we generate different subsequences of different combinations of static and dynamic tokens. In the selection stage, we evaluate their query-aware metrics based on shallow attention layers and select the optimal one for input to LLMs.
@@ 4.1 Problem Definition
Due to GPU memory and model capability constraints, the number of visual tokens processed during inference is capped at $L_{b}$. The fixed token budget limits frame sampling to a reduced number of frames, resulting in significant loss of rich visual information, particularly in long videos.
@@ 4.3 Exploration: Search Space Design
This section describes the generation of $n$ token subsequences, each of length $L_{b}$, from a token sequence of length $L$. To balance static and dynamic information in videos based on query requirements, we classify frames into static and dynamic frames. It is worth noting that since some models employ temporal reduction, here we refer to “frames” as $\bm{F}$ and the number of frames is $T$. Based on whether the tokens in a subsequence originate from static or dynamic frames, we divide them into two subsets: static frame tokens $\mathcal{T}_{\text{Static}}$ and dynamic delta tokens $\mathcal{T}_{\text{Dynamic}}$.
  • Static Frame Token. The static frame tokens are extracted from the static frames. Assuming $N_{s}$ static frames are chosen in the video, we uniformly select them from $\bm{F}$. The temporal indices of these frames are:
    =$\begin{gathered}\mathcal{I}=\left\{\left\lfloor\frac{kT}{N_{s}}\right\rfloor+1\mid k=0,1,\dots,N_{s}-1\right\},\end{gathered}$ (1)
    where the first frame is always selected as a static frame. We preserve all tokens from these static frames to constitute $\mathcal{T}_{\text{Static}}$ as:
@@
    where $\bm{F}^{i,h,w}\in\mathbb{R}^{D}$ represents the token embedding at the $i$-th frame and spatial location $(h,w)$ in $\bm{F}$. Obviously The total number of static tokens satisfies $|\mathcal{T}_{\text{Static}}|=N_{s}\times H\times W$.
  • Dynamic Delta Token. As illustrated in Figure 3, the static frames divide the entire sampled frame sequence into $N_{s}$ intervals. The frames within these intervals are defined as dynamic frames, and the dynamic delta tokens $\mathcal{T}_{\text{Dynamic}}$ are extracted from them to capture the dynamic information in the video. Since the subsequence length is $L_{b}$, it is evident that the number of dynamic delta tokens satisfies $|\mathcal{T}_{\text{Dynamic}}|=L_{b}-|\mathcal{T}_{\text{Static}}|$. These tokens are uniformly distributed across each interval, meaning the number of dynamic delta tokens selected from the $i$-th interval satisfy $|\mathcal{T}_{\text{Dynamic},i}|=\lfloor|\mathcal{T}_{\text{Dynamic}}|/N_{s}\rfloor$.
    Inspired by video codec, to retain as much dynamic information as possible, we select tokens from each interval that exhibit the largest differences compared to the corresponding tokens in the preceding static frame. We first define the token difference metric based on the cosine similarity between two token embeddings:
@@
    This metric increases as the two embeddings become more dissimilar. Then, we select $|\mathcal{T}_{\text{Dynamic},i}|$ tokens in the interval $i$ that have the largest differences compared to the corresponding tokens in the preceding static frame. We define $\mathcal{T}_{\text{Dynamic},i}$ as:
    $\begin{gathered}\{\bm{F}_{i}^{j,h,w}\mid\text{Top}_{|\mathcal{T}_{\text{Dynamic},i}|}\mathcal{D}(\bm{F}_{i}^{0,h,w},\bm{F}_{i}^{j,h,w}),\\ j\in[1,T_{i}],h\in[1,H],w\in[1,W]\},\end{gathered}$ (4)
    where $\bm{F}_{i}^{0,h,w}$ denotes the token embedding at position $(h,w)$ in the preceding static frame of current interval $i$, and $\bm{F}_{i}^{j,h,w}$ denotes the token embedding at position $(h,w)$ in the $j$-th dynamic frame of interval $i$. $T_{i}$ refers to the number of dynamic frames in $i$-th interval.
  • Token Subsequence Generation. Then we merge $\mathcal{T}_{\text{Static}}$ and $\mathcal{T}_{\text{Dynamic}}$ according to their original order to obtain the $L_{b}$-long token subsequence $\bm{\hat{T}}_{v}$.
To generate $n$ candidate token subsequences, we vary $N_{s}$ from $1$ to $n$. As $N_{s}$ decreases, the number of dynamic delta tokens increases, thereby preserving more dynamic information under the same token budget. Conversely, the number of static frame tokens increases, preserving more static information. In this way, we can generate token subsequences with varying proportions of static and dynamic information to adapt to the requirements of different queries.
Figure 3: An example of token subsequence generation with 6 total frames and 2 static frames.
@@ 4.4 Selection: Quick Evaluation
After obtaining $n$ token subsequences of length $L_{b}$, we perform an evaluation and select the optimal subsequence based on the chosen metric.
Previous studies have identified certain characteristics of visual tokens in attention mechanisms. For instance, Chen et al. (2024a) shows that most vision tokens can be removed at the second layer without significant performance loss, and Wan et al. (2024) observes that vision tokens are generally less attended. Based on these findings, we assume that the attention mechanism at the second layer already provides meaningful clues of token importance. Besides, we hypothesize that inputs with more tokens being attended are more likely to yield better outputs.
To enable quick evaluation, we compute the attention score matrix $\bm{S}$ at the second layer of the VideoLMs, using textual query tokens as the query input, and instruction and vision tokens as the key input:
@@
$\bm{K}=\bm{W}_{\text{K}}\ \cdot\text{Concat}(\bm{H}_{i},\bm{H}_{v}),$ (6)
@@
where $\bm{H}_{q}$, $\bm{H}_{i}$, and $\bm{H}_{v}$ denote the hidden features of the textual query, instruction, and visual inputs. And $d_{k}$ is the dimension of key vectors in the attention mechanism. To quantify the attention allocated to visual tokens, we compute the summation of attention scores of the visual tokens. Specifically, to ensure comprehensive consideration of each text query token, we first extract the maximum values along the query dimension from $\bm{S}$, yielding an attention score vector $\bm{s}$ for each visual token. Then we sum the attention scores of the visual tokens:
@@
${\bm{\bar{T}}_{v}}=\underset{m\in\{1,2,\dots,n\}}{\arg\max}s^{(m)},$ (10)
where $s^{(m)}$ denotes the summed attention score for the $m$-th token subsequence.

### v2

@@ Abstract
Video question answering benefits from the rich information in videos, enabling various applications. However, the large volume of tokens generated from long videos presents challenges to memory efficiency and model performance. To alleviate this, existing works propose to compress video inputs, but often overlook the varying importance of static and dynamic information across different queries, leading to inefficient token usage within limited budgets. We propose a novel token selection strategy, explore-then-select, that adaptively adjusts static and dynamic information based on question requirements. Our framework first explores different token allocations between key frames, which preserve spatial details, and delta frames, which capture temporal changes. Then it employs a query-aware attention-based metric to select the optimal token combination without model updates. Our framework is plug-and-play and can be seamlessly integrated within diverse video language models. Extensive experiments show that our method achieves significant performance improvements (up to 5.8%) on multiple video question answering benchmarks. Our code is available at https://github.com/ANDgate99/Explore-Then-Select.
Static or Dynamic: Towards Query-Adaptive Token Selection for
Video Question Answering
@@ 1 Introduction
Video Question Answering (VideoQA) has broad applications across various fields Mogrovejo and Solorio (2024); Zhang et al. (2024a). Compared to text, videos provide more intuitive and dynamic information, delivering richer context and details by combining visual and temporal elements. Current research primarily leverages powerful large language models to build video language models (VideoLMs) Lin et al. (2023); Zhang et al. (2024b), significantly enhancing AI performance in VideoQA tasks. However, the extensive visual information in long videos leads to a dramatic increase in token counts. For instance, if one frame generates 196 tokens Li et al. (2024a), a 5-minute video sampled at 1 fps would produce nearly 60,000 tokens, posing significant challenges to memory requirements and model capabilities.
Figure 1: Different question types vary in their dependence on static and dynamic information in videos. For example, Question 2 relies on fine-grained dynamic information, while Question 1 and 3 only require key frames. The frames needed to answer the questions are highlighted with corresponding colored boxes.
@@
To achieve an effective allocation between static and dynamic information in visual token compression, we propose a novel token selection strategy, Explore-then-Select, that adaptively aligns visual tokens with textual queries under a limited token budget. Unlike previous approaches that rely on fixed rules, our strategy autonomously and adaptively combines static and dynamic content based on the nature of the questions (e.g., action description, event sequence, or object recognition), ensuring more precise responses to diverse queries.
Specifically, we categorize video frames into key and delta frames. Key frames are fully retained to preserve essential spatial details, such as objects, while delta frames are sparsely processed, keeping only a subset of tokens to capture important temporal changes. To optimize token allocation between these two types of frames, Explore-then-Select uses a two-stage process. In the exploration stage, we construct a search space comprising various combinations of key and delta frames, each yielding a token subsequence of constrained length. By adjusting the proportion of key and delta frames, we can prioritize either static details or dynamic changes based on the question requirements. In the selection stage, we evaluate each combination using a query-aware metric derived from the shallow attention layers of VideoLMs. This metric quantifies the alignment between the query and visual tokens, enabling us to select the optimal combination to answer the question.
@@
  • Building on the observation that questions rely differently on static and dynamic video information, we propose a novel Explore-then-Select framework to adaptively and effectively select visual tokens reflecting the optimal balance of static and dynamic information under limited token budgets.
  • To address static and dynamic information needs, we design an effective search space of key-delta frame combinations. During the selection phase, we employ a query-aware approach, leveraging an attention-based metric to adaptively evaluate candidates and select the optimal combination for each question.
  • We conduct extensive experiments on both long and short video benchmarks, demonstrating the effectiveness of our method. Thanks to its plug-and-play design, our approach generalizes well across different models without extra fine-tuning and enables direct control over the token budget for flexible adaptation to resource constraints.
@@ 2.1 Video Language Models
Significant progress has been made in video language model research based on LLMs. These models can be primarily classified into two types: general-purpose vision language models Team et al. (2024); Chen et al. (2024b); OpenAI (2024); Yao et al. (2024); Ye et al. (2023) and specialized video language models Lin et al. (2023); Zhang et al. (2024c); Li et al. (2024c); Zhang et al. (2025); Liu et al. (2024). Among the former, LLaVA-OneVision Li et al. (2024a) unifies image and video tasks, while Qwen2-VL Wang et al. (2024) introduces dynamic resolution support and three-dimensional positional encoding for enhanced visual feature capture. Among specialized models, VideoChat Li et al. (2023b) targets deep video understanding and interaction, and LongVA Zhang et al. (2024b) extends the context length of language models, transferring their advantages in long-text processing to the video domain.
@@
FastV Chen et al. (2024a)
ZipVL He et al. (2024b)
@@
[…]
Table 1: Feature comparison with existing methods. “Pre-Input” refers to methods that reduce tokens before feeding them into large language models, while “Video-Specific” denotes methods that leverage the unique characteristics of video data.
@@ 2.2 Visual Token Compression
Some studies Bolya et al. (2022) focus on compressing visual tokens in vision encoders. For example, RLT Choudhury et al. (2024) effectively reduces the number of tokens by replacing repeated patches in videos with a single patch. Other works Li et al. (2024b); Qian et al. (2025); Shen et al. (2024); Lan et al. (2024) introduce dedicated modules for token compression, such as BLIP-2 Li et al. (2023a), which uses a Q-Former module with learnable queries to generate compact semantic representations. Additionally, inspired by KV cache compression in long text processing Zhang et al. (2023), some methods apply similar strategies to visual tokens He et al. (2024b); Chen et al. (2024a); Fu et al. (2024b). These methods optimize token usage efficiency by setting thresholds based on specific metrics to prune visual tokens.
Table 1 compares existing methods, noting that training-free approaches mainly compress tokens within the KV cache, reducing FLOPs but failing to address the issue of excessive token input to large language models. In contrast, methods that reduce tokens in advance typically require training. This paper introduces a novel pre-input, training-free framework for more effective compression, balancing query-aware static and dynamic information.
@@ 3 Preliminary
  • Visual Encoding and Embedding. The sampled frames are divided into non-overlapping spatiotemporal patches, which are processed by a vision encoder to extract spatiotemporal features. These features are then projected into the language model’s token space via a linear projection, resulting in visual token embeddings $\bm{F}\in\mathbb{R}^{T\times H\times W\times D}$, where $T$ represents the temporal resolution (typically equal to $N$ unless temporal downsampling is applied), $H$ and $W$ denote the spatial resolutions, and $D$ is the token embedding dimension.
  • Multimodal Processing. The visual token embeddings $\bm{F}$ are then flattened into a sequence $\bm{T}_{v}\in\mathbb{R}^{L\times D}$, where $L=T\times H\times W$ is the sequence length. The sequence $\bm{T}_{v}$, instruction embeddings $\bm{T}_{i}$, and query embeddings $\bm{T}_{q}$ are concatenated into a unified input $\bm{T}=[\bm{T}_{i},\bm{T}_{v},\bm{T}_{q}]$, where $\cdot$] denotes token concatenation. Finally, the VideoLM processes the unified input sequence $\bm{T}$ to generate a textual response to the question.
Figure 2: Overview of our Explore-then-Select framework for token selection. During the exploration stage, multiple subsequences are generated from different combinations of key and delta-frame tokens. In the selection stage, these subsequences are evaluated using query-aware metrics computed from shallow attention layers, and the optimal subsequence is chosen as input to the LLM.
@@ 4.1 Problem Definition
Due to GPU memory and model capability constraints, the number of visual tokens processed during inference is capped at $L_{b}$. The fixed token budget limits frame sampling to a reduced number of frames, resulting in a significant loss of rich visual information, particularly in long videos.
@@ 4.3 Exploration: Search Space Design
In this section, we describe the generation of $n$ token subsequences, each of length $L_{b}$, from a token sequence of length $L$. To balance static and dynamic information in videos according to query requirements, we classify frames into key and delta frames. Note that, due to temporal downsampling in some models, “frames” here refer to visual token embeddings $\bm{F}$, and the total number of frames is $T$. Based on whether the tokens in a subsequence originate from key or delta frames, we divide them into two subsets: key-frame tokens $\mathcal{T}_{\text{key}}$ and delta-frame tokens $\mathcal{T}_{\text{delta}}$.
  • Key-frame Token. The key-frame tokens are extracted from the key frames. Assuming $N_{s}$ key frames are selected in the video, we select them uniformly from $\bm{F}$. The temporal indices of these frames are:
    $\begin{gathered}\mathcal{I}=\left\{\left\lfloor\frac{kT}{N_{s}}\right\rfloor+1\mid k=0,1,\dots,N_{s}-1\right\},\end{gathered}$ (1)
    where the first frame is always selected as a key frame. All tokens from these key frames are retained to form $\mathcal{T}_{\text{key}}$:
@@
    where $\bm{F}^{i,h,w}\in\mathbb{R}^{D}$ represents the token embedding at the $i$-th frame and spatial location $(h,w)$ in $\bm{F}$. Hence, the total number of key-frame tokens is $|\mathcal{T}_{\text{key}}|=N_{s}\times H\times W$.
  • Delta-frame Token. As illustrated in Figure 3, the key frames partition the entire sampled frame sequence into $N_{s}$ intervals. The frames between consecutive key frames within each interval are defined as delta frames, and delta-frame tokens $\mathcal{T}_{\text{delta}}$ are extracted from them to capture dynamic information relative to the preceding key frames. Given a subsequence length of $L_{b}$, the total number of delta-frame tokens is $|\mathcal{T}_{\text{delta}}|=L_{b}-|\mathcal{T}_{\text{key}}|$. These tokens are uniformly distributed across the intervals, such that the number of delta-frame tokens selected from the $i$-th interval is $|\mathcal{T}_{\text{delta},i}|=\lfloor|\mathcal{T}_{\text{delta}}|/N_{s}\rfloor$.
    Inspired by video coding techniques, to retain as much dynamic information as possible, we select tokens from each interval that exhibit the largest differences compared to the corresponding tokens in the preceding key frame. We first define the token difference metric based on the cosine similarity between two token embeddings:
@@
    This metric increases as the two embeddings become more dissimilar.
    For interval $i$, we compute the difference between each token in the frames within the interval and the token at the corresponding spatial position in the preceding key frame. Specifically, we define:
    $\begin{gathered}\Delta_{i}(j,h,w)=\mathcal{D}(\bm{F}_{i}^{0,h,w},\bm{F}_{i}^{j,h,w})\\ j\in[1,T_{i}],h\in[1,H],w\in[1,W],\end{gathered}$ (4)
    where $\bm{F}_{i}^{0,h,w}$ denotes the token embedding at spatial location $(h,w)$ in the preceding key frame of interval $i$, and $\bm{F}_{i}^{j,h,w}$ denotes the token embedding at the same spatial location in the $j$-th delta frame within the interval. Here, $T_{i}$ is the number of delta frames contained in interval $i$.
    We then select the delta-frame tokens corresponding to the $|\mathcal{T}_{\text{delta},i}|$ largest values of $\Delta_{i}(j,h,w)$, forming the set $\mathcal{T}_{\text{delta},i}$ as:
    $\{\bm{F}_{i}^{j,h,w}\mid(j,h,w)\in\operatorname{arg\,Top}_{|\mathcal{T}_{\text{delta},i}|}\Delta_{i}(j,h,w)\}.$ (5)
  • Token Subsequence Generation. Then we merge $\mathcal{T}_{\text{key}}$ and $\mathcal{T}_{\text{delta}}$ according to their original order to obtain the $L_{b}$-long token subsequence $\bm{\hat{T}}_{v}$.
Figure 3: An example of token subsequence generation with 6 total frames and 2 key frames.
To generate $n$ candidate token subsequences, we vary the number of key frames $N_{s}$ from $1$ to $n$. A smaller $N_{s}$ results in more delta-frame tokens, thereby capturing more dynamic information within the same token budget. Conversely, a larger $N_{s}$ increases the number of key-frame tokens, preserving more static information. In this way, we can generate token subsequences with varying proportions of static and dynamic information to adapt to the requirements of different queries.
Notably, our frame division is inspired by the GOP structure in video codec Lee et al. (2006), where I-frames encode full scene information and P/B-frames encode temporal differences. Similar to adjusting GOP sizes, varying the proportion of key and delta frames allows us to control the emphasis on static or dynamic cues.
@@ 4.4 Selection: Quick Evaluation
After obtaining $n$ token subsequences of length $L_{b}$, we perform an evaluation and select the optimal subsequence based on the chosen metric. Previous studies have identified certain characteristics of visual tokens in attention mechanisms. For instance, Chen et al. (2024a) shows that most visual tokens can be removed at the second layer without significant performance loss, and Wan et al. (2024) observes that visual tokens are generally less attended. Based on these findings, we consider that the attention mechanism at the second layer already provides meaningful clues of token importance. Besides, we hypothesize that higher cumulative attention scores on visual tokens indicate a better utilization of the visual information.
Model Settings EgoSchema VideoMME MLVU
Method Sample Budget Short Medium Long Overall
VideoChat2 - 16 - 54.4 48.3 37.0 33.2 39.5 -
LongVA - 128 - - 61.1 50.4 46.2 52.6 -
mPLUG-Owl3 - 128 - - 70.0 57.7 50.1 59.3 -
LongVU - 1fps - 67.6 - - - 60.6 65.4
Qwen2-VL-7B Original 64 - 66.2 71.1 59.4 50.8 60.4 50.6
Retrieval 256 64 63.6 71.0 61.3 52.2 61.5 49.4
Similarity 256 64 66.6 71.4 60.6 51.8 61.3 53.0
[…]
Original 32 - 64.7 68.9 55.2 48.7 57.6 46.8
Retrieval 128 32 61.7 70.0 58.6 51.6 60.0 46.8
Similarity 128 32 65.6 70.1 58.7 51.8 60.2 47.2
[…]
LLaVA-OneVision-7B Original 64 - 60.1 70.6 55.8 47.8 58.0 50.8
Retrieval 256 64 57.7 64.0 53.4 47.0 54.8 44.6
Similarity 256 64 59.6 71.0 57.9 50.8 59.9 48.4
[…]
Original 32 - 60.4 71.3 57.4 48.0 58.9 46.8
Retrieval 128 32 57.9 63.2 53.9 46.0 54.4 44.0
Similarity 128 32 60.2 70.8 57.1 49.7 59.2 50.2
[…]
Table 2: Results on long video benchmarks show that our method achieves significant improvements over the baselines, particularly on the advanced Qwen2-VL, with up to a 5.8% gain on the VideoMME medium subset.
To enable quick evaluation, we compute the attention score matrix $\bm{S}$ at the second layer of the VideoLMs, using textual query tokens as the query input, and instruction and visual tokens as the key input:
@@
$\bm{K}=\bm{W}_{\text{K}}\,\text{concat}(\bm{H}_{i},\bm{H}_{v}),$ (7)
@@
where $\bm{H}_{q}$, $\bm{H}_{i}$, and $\bm{H}_{v}$ denote the hidden features of the textual query, instruction, and visual inputs. Here, $d_{k}$ is the dimension of key vectors in the attention mechanism. To quantify the attention allocated to the visual tokens, we compute the summed attention scores of the visual tokens. Specifically, to ensure comprehensive consideration of each text query token, we first extract the maximum values along the query dimension from $\bm{S}$, yielding an attention score vector $\bm{s}$ for each visual token. Then we sum the attention scores of the visual tokens:
@@
${\bm{\bar{T}}_{v}}=\underset{m\in\{1,2,\dots,n\}}{\arg\max}s^{m},$ (11)
where $s^{m}$ denotes the summed attention score for the $m$-th token subsequence.